Tina rides her bike 15m up a hill and then 15m. down the same hill. This complete trip takes 2 hours. If her downhill speed is 20m faster than her uphill speed, how fast does she ride uphill?
I have these variables and have tried them every which way but loose...
uphill- distance=15, rate=r, and time=t ;
downhill- distance=15, rate=r+20, and time=2-t
No matter what I do, nothing looks workable for me...I think my brain is stuck.
Thank you, BP

Answer by Alan3354(69443)

Tina rides her bike 15m up a hill and then 15m. down the same hill. This complete trip takes 2 hours. If her downhill speed is 20m faster than her uphill speed, how fast does she ride uphill?
---------
I'll assume you mean 20 km/hr faster, and that the distances are km, not meters.
----------
d = rt (r = uphill rate)
t = d/r
2 hr = 15/r + 15/(r+20)
2*r*(r+20) = 15(r+20) + 15r
2r^2 + 40r = 15r+300 + 15r
2r^2 + 10r - 300 = 0
r^2 + 5r - 150 = 0
(r + 15)(r - 10) = 0
r = 10 km/hr uphill, 30 km/hr downhill